Nuc 1 is fixed?
Swarm still up in close by tree. No luck at shaking them down. Cluster looks to be about 3-4 lbs of bees.
Nuc#1 Frame of eggs moved from Nuc#2 are capped with 3 queen cells created. No signs if a laying queen.
Nuc#2 Queen located. Bright orange with a bit of black marking at the tip. She was not moving very fast, but there was eggs and larva found at various stages.
Nuc#3 Located at other property, pollen being brought in, new foragers doing orientation flights.
Warre#1 Swarmed yesterday. Did not inspect to avoid damaging queen cells.
Warre#2 Nice drawn out comb with brood at various stages. New foragers expected at the end of this month.
hTBH Top lang super is still being drawn out. Empty frames have bees festooning, but are not being drawn out. Looked at on bar to note queen cells. Queens are expected to emerge ~1 week from tomorrow.